The Department of Computer Science and Engineering

University of Nevada, Reno

cordially invites you to a Master’s colloquium

Ringermute: An audio data mining toolkit

A thesis submitted in partial fulfillment of the requirements for the degree of Master of Science with a major in Computer Science.

by

Marcel Andrew Levy

Abstract: This thesis presents Ringermute, an application designed to support audio feature recognition and machine learning, from the training and testing to the deployment phase. By choosing from a combination of feature extraction routines provided by plug-ins, a researcher can quickly produce files for input to standard data mining tools. The best combination of featureextraction and classifier plugins can then be used to drive a near-real-time application for further testing or production use.

10:00 am, Tuesday, November 22, 2005

Access Grid Node, Scrugham Engineering and Mines (SEM) room 201

for more information contact Dr. Fred Harris @ 784-6571 (fredh@cse.unr.edu)